OpenStudy (dlordy132):

Identify the domain and range of the function y=-0.25x^2

OpenStudy (anonymous):

you can square anything

OpenStudy (dlordy132):

?

OpenStudy (anonymous):

you know what "domain" means? "no" is a fine answer, just asking

OpenStudy (dlordy132):

not really :P

OpenStudy (anonymous):

possible values of x, the input

OpenStudy (anonymous):

so when i said "you can square any number" that means the domain is assumed to be all real numbers

OpenStudy (dlordy132):

ohh what about the range?

OpenStudy (anonymous):

that is the possible outputs

OpenStudy (anonymous):

since \(x^2\) is never negative , you know \(x^2\geq 0\) always that means \(-x^2\leq 0\) and that is your range \(y\leq 0\)

OpenStudy (dlordy132):

oh ok thank you

OpenStudy (anonymous):

yw